An office heater puts out 2050 Btus of energy in the form of heat per hour. Given that 1 Btu=1.055kJ how many?
megajoules of energy in the form of heat can be produced per year by the heater
1
Expert's answer
2017-09-04T09:05:06-0400
2050Btu-in1 Hour
1Btu=1.055Kj
Q in year
1 year=365 days=8760 hours
Q(energy)=8760*2050=17958000Btu=17959000*1.055=18945.69MJ